OpenStudy (anonymous):

I need the solutn of dis progm...plz help... write a prgm in C or C++ 4r swaping 2 variable's value without using any temporary variable...

OpenStudy (anonymous):

use array

OpenStudy (anonymous):

@thangeram can u post d prgm here...

OpenStudy (anonymous):

sry my array method is not giving me satisfying output bt u can try this a = a + b; b = a - b; a = a - b;

OpenStudy (anonymous):

If they hold different (integral) values, use the following swap. template<typename T> inline void swap(T &l, T &r) { l ^= r; r ^= l; l ^= r; } Why don't you just use std::swap?

OpenStudy (anonymous):

You need to be very careful using the methods above. You may introduce a bug that causes some values to work and others to fail (this is due to overflow). I believe xor is safer than addition in this case.

OpenStudy (anonymous):

thanks for reply...

OpenStudy (anonymous):

The XOR will not cause overflow.